### v1

# CALT-TH 2023-050

Abstract
We show via explicit construction that all the extreme rays of both the 3-party quantum entropy cone and the 4-party stabilizer entropy cone can be obtained from subsystem coarse-grainings of specific higher-party quantum states, namely extreme states characterized by saturating a (non-trivial) maximal set of instances of subadditivity. This suggests that the study of the “subadditivity cone”, and the set of its extreme rays realizable in quantum mechanics, provides a powerful new tool for deriving inner bounds for the quantum and stabilizer entropy cones, as well as constraints on new inequalities for the von Neumann entropy.
I Introduction
The derivation of the fundamental inequalities satisfied by the von Neumann entropy is an important problem in quantum information theory. Unfortunately, it is also notoriously difficult. Since the proof of strong subadditivity by Lieb and Ruskai half a century ago Lieb and Ruskai (1973a, b), no new unconstrained inequality has been found. New inequalities involving four or more parties were discovered in Linden and Winter (2005); Cadney et al. (2012), but they are constrained, meaning that they only apply to particular density matrices which saturate other inequalities. In the classical case, i.e., for the Shannon entropy, new inequalities have been derived in Zhang and Yeung (1998); Makarychev et al. (2002); Zhang (2003); Matús (2007), and Cadney et al. (2012) speculated that they might hold also in quantum mechanics. However, to our knowledge, at present there is no strong argument that new unconstrained inequalities for the von Neumann entropy should exist.
The main goal of this letter is to show that a seemingly much simpler problem, the quantum marginal independence problem (QMIP) Hernández-Cuenca et al. (2019), originally motivated by an approach Hubeny et al. (2018, 2019); Hernández-Cuenca et al. (2022) to the study of entropy inequalities in quantum gravity Ryu and Takayanagi (2006); Hayden et al. (2013); Bao et al. (2015), might provide a powerful tool for the derivation of constraints on the existence of new inequalities and their structure. Specifically, we will show that already from a partial solution to the “extremal” version of the QMIP up to nine parties, we can derive definite inner bounds for the quantum entropy cone, which in the case of three parties coincide with the full cone Pippenger (2003), and in the case of four parties coincide with the entropy cone of stabilizer states Linden et al. (2013).
The structure of the paper is as follows. In §II and §III, we review the main definitions concerning entropy cones, the QMIP, and its extremal version. In §IV, we formulate the inner bound to the quantum entropy cone for an arbitrary number of parties. In §V, we review the definition of hypergraph models from Bao et al. (2020a) and the main result from Walter and Witteveen (2021) about their realizability by stabilizer states. We then use this technology in §VI to derive the main result of this letter, and conclude in §VII with a list of open questions.
II The quantum entropy cone
A convenient framework to study entropy inequalities (and their relations) was introduced in Pippenger (2003), following the analogous work of Yeung (1997) for the Shannon entropy. Let $[{\sf N}]=\{1,2,\ldots,{\sf N}\}$, and consider an ${\sf N}$-party density matrix $\rho_{{\sf N}}$ on a Hilbert space $\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\ldots\otimes\mathcal{H}_{{\sf N}}$. The entropy vector of $\rho_{{\sf N}}$ is the vector in $\mathbb{R}^{\sf{D}}$, with $\sf{D}=2^{{\sf N}}-1$, given by
$\vec{{\sf S}}=\{{\sf S}_{\mathscr{I}},\;\forall\,\mathscr{I}\subseteq[{\sf N}]\},\qquad{\sf S}_{\mathscr{I}}:={\sf S}(\rho_{\mathscr{I}}),$ (1)
where $\rho_{\mathscr{I}}$ is the reduced density matrix for the parties in $\mathscr{I}$. Denoting by $\Sigma_{{\sf N}}^{*}$ the set of entropy vectors for all ${\sf N}$-party quantum states, its topological closure $\overline{\Sigma}^{*}_{{\sf N}}$ is a convex cone Pippenger (2003) known as the quantum entropy cone (QEC).
For any ${\sf N}$, an outer bound of $\overline{\Sigma}^{*}_{{\sf N}}$ is given by the polyhedral cone obtained via the intersection of the half-spaces specified by all instances of subadditivity (SA) and strong subadditivity (SSA) of the von Neumann entropy, and we will denote this cone by $\Sigma_{{\sf N}}$. For ${\sf N}\leq 3$ it is easy to show that $\overline{\Sigma}^{*}_{{\sf N}}=\Sigma_{{\sf N}}$ by constructing quantum states that realize entropy vectors belonging to the extreme rays of $\Sigma_{{\sf N}}$. On the other hand, for ${\sf N}\geq 4$, the constrained inequalities found in Linden and Winter (2005); Cadney et al. (2012) imply that this construction is no longer possible, since there are regions on the boundary of $\Sigma_{{\sf N}}$ which are inaccessible to quantum states. This however is not enough to conclude that $\overline{\Sigma}^{*}_{{\sf N}}\subset\Sigma_{{\sf N}}$ (strictly), since there might be quantum states that approximate the entropy vectors in these regions arbitrarily well.
III Quantum marginal independence
An interesting and well-known problem in quantum mechanics, where entropy inequalities play a central role, is the quantum marginal problem Schilling (2014). Given density matrices $\rho_{\mathscr{I}}$ for some subsystems, it asks whether there exists a global density matrix $\rho$ such that all the given $\rho_{\mathscr{I}}$ can be obtained from $\rho$ as marginals. Entropy inequalities restrict the set-ups for which a solution exists.
The quantum marginal independence problem (QMIP) Hernández-Cuenca et al. (2019) can be interpreted as a simplified version of the quantum marginal problem, where instead of fixing the marginals for some subsystems, one only demands that certain subsystems are correlated and others are not. Specifically, the QMIP asks the following question: Given an ${\sf N}$-party system and a complete specification of the presence of correlation (or conversely, the lack thereof) among the various subsystems, is there a density matrix that satisfies these constraints? As for the case of the marginal problem, entropy inequalities can be used to constrain the space of solutions for the QMIP. However, unlike its parent version, the knowledge of the quantum entropy cone is sufficient to solve the problem completely. To see this, let us formulate the QMIP more precisely.
The ${\sf N}$-party subadditivity cone (SAC) is defined as the polyhedral cone in entropy space carved out by all instances of SA for that ${\sf N}$.
Consider a face $\mathscr{F}$ of the SAC and a vector $\vec{\sf S}\in\text{int}(\mathscr{F})$. Notice that the collection of SA instances which are saturated by $\vec{\sf S}$ is independent from the specific choice of this vector. The saturation of SA is equivalent to the vanishing of the mutual information, which occurs if and only two subsystems are independent. We can then interpret $\text{int}(\mathscr{F})$ to correspond to a specification of which subsystems are correlated and which are independent, while remaining agnostic about the specific values of the entropies. In its original formulation Hernández-Cuenca et al. (2019), the QMIP asked for which faces does $\text{int}(\mathscr{F})$ contain at least one entropy vector that can be realized by a quantum state. We now slightly generalize this question by replacing “realized” with “approximated arbitrarily well.” In this generalized version, the QMIP can then be trivially solved if $\overline{\Sigma}^{*}_{{\sf N}}$ is known.
In this letter we proceed in the opposite direction and attempt to extract information about $\overline{\Sigma}^{*}_{{\sf N}}$ from the solution to the QMIP for different values of ${\sf N}^{\prime}\geq{\sf N}$. In particular, we will be interested in the extremal version of the QMIP, which we denote as EQMIP, where we only focus on the 1-dimensional faces of the SAC, i.e., its extreme rays. Denoting by ^$\widehat{\mathscr{R}}_{{\sf N}}$ the set of all extreme rays of the ${\sf N}$-party SAC, the solution $\mathscr{R}_{{\sf N}}$ to the EQMIP is then
$\mathscr{R}_{{\sf N}}:=\widehat{\mathscr{R}}_{{\sf N}}\cap\overline{\Sigma}^{*}_{{\sf N}}$ (2)
We will now explain how to construct inner bounds for $\overline{\Sigma}^{*}_{{\sf N}}$ from the knowledge of $\mathscr{R}_{{\sf N}^{\prime}}$ for some ${\sf N}^{\prime}\geq{\sf N}$.
IV Inner bounds from extremal marginal independence
IV.1 Subsystem coarse-grainings
Given a density matrix $\rho_{{\sf N}^{\prime}}$, we want to consider a coarse-graining of the ${\sf N}^{\prime}$ parties into ${\sf N}$ composite ones and to relate the entropy vectors before and after the coarse-graining. We will also consider purifications of $\rho_{{\sf N}^{\prime}}$, and allow for coarse-grainings of the “purifying” party. Defining $\llbracket{\sf N}\rrbracket=\{0,1,\ldots,{\sf N}\}$, where we added 00 to account for the purifier, such a coarse-graining can then be specified by a surjective function
$\begin{split}f:\;&\llbracket{\sf N}^{\prime}\rrbracket\rightarrow\;\llbracket{\sf N}\rrbracket\\ &\ell^{\prime}\;\mapsto\;\ell=f(\ell^{\prime}),\end{split}$ (3)
which specifies, for each party $\ell^{\prime}\in\llbracket{\sf N}^{\prime}\rrbracket$, which of the coarse-grained $\llbracket{\sf N}\rrbracket$ parties it belongs to.
Given a density matrix $\rho_{{\sf N}^{\prime}}$ with entropy vector $\vec{{\sf S}}^{\prime}$, and a coarse-graining $f$, the components of the entropy vector of the same density matrix after the coarse graining are then given by
${\sf S}_{\mathscr{I}}={\sf S}_{f^{-1}(\mathscr{I})},$ (4)
where $f^{-1}(\mathscr{I})$ is the pre-image of $\mathscr{I}$  [^1] If $0\in f^{-1}(\mathscr{I})$, then $f^{-1}(\mathscr{I})\not\subseteq[{\sf N}^{\prime}]$ and ${\sf S}_{f^{-1}(\mathscr{I})}$ is not a component of $\vec{{\sf S}}(\rho_{{\sf N}^{\prime}})$. However, the entropy of its complement in $\llbracket{\sf N}^{\prime}\rrbracket$ is and is equal to ${\sf S}_{f^{-1}(\mathscr{I})}$. This is what the right-hand side of Eq. (4) refers to implicitly..
Notice that in general, this map of entropy vectors from $\Sigma_{{\sf N}^{\prime}}^{*}$ to $\Sigma_{{\sf N}}^{*}$, which we denote by $\Psi_{f}$, can formally be extended to vectors in the full $\overline{\Sigma}^{*}_{{\sf N}^{\prime}}$, and can map a vector on the boundary of $\overline{\Sigma}^{*}_{{\sf N}^{\prime}}$ to one strictly in the interior of $\overline{\Sigma}^{*}_{{\sf N}}$. For example, any coarse-graining to ${\sf N}=2$ of the ${\sf N}=3$ entropy vector corresponding to the 4-party GHZ state gives an entropy vector strictly inside $\overline{\Sigma}^{*}_{2}$ (since no instance of SA is saturated).
IV.2 Inner bounds for $\overline{\Sigma}^{*}_{{\sf N}}$
Using this construction, we can now derive an inner bound for $\overline{\Sigma}^{*}_{{\sf N}}$ as follows. Suppose that $\mathscr{R}_{{\sf N}^{\prime}}$ is known for some ${\sf N}^{\prime}\geq{\sf N}$. We can then consider all possible coarse-graining from ${\sf N}^{\prime}$ to ${\sf N}$, mapping all vectors in $\mathscr{R}_{{\sf N}^{\prime}}$ to ${\sf N}$-party entropy vectors, and taking the conical hull. Formally, for any given ${\sf N}$ and ${\sf N}^{\prime}\geq{\sf N}$, we define
$\Delta_{{\sf N}}^{{\sf N}^{\prime}}=\text{cone}\,\big{\{}\Psi_{f}(\vec{{\sf R}}),\;\forall f\in\Phi\;\;\forall\,\vec{\sf R}\in\mathscr{R}_{{\sf N}^{\prime}}\big{\}},$ (5)
where $\Phi$ is the set of all possible functions $f$ from Eq. (3). We then have the following lemma.
Lemma 1.
For any fixed ${\sf N}$, and any $i,j\in\mathbb{N}$ with $i<j$, $\Delta_{{\sf N}}^{{\sf N}+i}\subseteq\Delta_{{\sf N}}^{{\sf N}+j}$.
Proof.
It suffices to show that for any ${\sf N}^{\prime}\geq{\sf N}$ and $\vec{\sf R}\in\mathscr{R}_{{\sf N}}$, there exists some $\vec{\sf R}^{\prime}\in\mathscr{R}_{{\sf N}^{\prime}}$ and a coarse-graining $f$ from ${\sf N}^{\prime}$ to ${\sf N}$ such that $\Psi_{f}(\vec{\sf R}^{\prime})=\vec{\sf R}$. For the case where $\vec{\sf R}$ is realized by a density matrix $\rho$, we can realize the desired $\vec{\sf R}^{\prime}$ with $\rho^{\prime}=\rho\otimes\ket{{\bf 0}}\!\bra{{\bf 0}}_{[{\sf N}^{\prime}]\setminus[{\sf N}]}$, and choose $f$ such that the additional factors are coarse-grained with the purifier for $\rho^{\prime}$. For the general case, we refer the reader to Section 3.4 of He et al. (2023). ∎
By Lemma 1 and the definition of $\Delta_{{\sf N}}^{{\sf N}^{\prime}}$ in Eq. (5), we obtain the following chain of inclusions:
$\Delta_{{\sf N}}^{{\sf N}}\subseteq\Delta_{{\sf N}}^{{\sf N}+1}\subseteq\Delta_{{\sf N}}^{{\sf N}+2}\subseteq\ldots\subseteq\overline{\Sigma}^{*}_{{\sf N}}\subseteq\Sigma_{{\sf N}}.$ (6)
It is then natural to ask how well this sequence approximates $\overline{\Sigma}^{*}_{{\sf N}}$, and if there is some “maximal” $\widehat{i}$ (which implicitly depends on ${\sf N}$) such that
$\Delta_{{\sf N}}^{{\sf N}+\widehat{i}}=\Delta_{{\sf N}}^{{\sf N}+i},\quad\forall\,i>\widehat{i}.$ (7)
In the following sections, we will answer this question for ${\sf N}=3$ and provide a partial answer for ${\sf N}=4$. First, however, we need to review a construction that will allow us to determine at least a subset of solutions to the EQMIP for sufficiently large ${\sf N}$.
V Hypergraph models
The hypergraph models of entanglement were introduced in Bao et al. (2020a) as a generalization of the graph models used in Bao et al. (2015) to study entropy inequalities in quantum gravity. An ${\sf N}$-party hypergraph model is a simple  [^2] By simple we mean the hypergraph has no loops and no repeated edges. weighted hypergraph $H=(V,E)$ with vertices $V$, hyperedges $E$ with positive weights, a specification of a subset $\partial V\subseteq V$ of vertices called boundary vertices, and a surjective (but not necessarily injective) map $\xi:\partial V\rightarrow\llbracket{\sf N}\rrbracket$.
Given an ${\sf N}$-party hypergraph model, one associates to it an entropy vector as follows. For a non-empty subset $\mathscr{I}\subseteq[{\sf N}]$, an $\mathscr{I}$-cut is a subset $V_{\mathscr{I}}\subset V$ such that $\partial V\cap V_{\mathscr{I}}=\xi^{-1}(\mathscr{I})$, where $\xi^{-1}$ denotes the pre-image. The cost of any such cut is the sum of the weights of the hyperedges that connect a vertex in $V_{\mathscr{I}}$ to one in $V_{\mathscr{I}}^{c}$, the complement of $V_{\mathscr{I}}$ in $V$. In other words, given an $\mathscr{I}$-cut $V_{\mathscr{I}}$, and a hyperedge $h$ (thought of as a collection of vertices), the weight of $h$ contributes to the cost of the cut if and only if $h$ contains at least one vertex in $V_{\mathscr{I}}$ and at least one in $V_{\mathscr{I}}^{c}$. The entropy ${\sf S}_{\mathscr{I}}$ is then defined as the cost of the $\mathscr{I}$-cut with minimal cost.
The prescription we just described associates to each ${\sf N}$-party hypergraph model a vector in $\mathbb{R}^{\sf{D}}$. The collection of all these vectors (at fixed ${\sf N}$) is again a convex cone  [^3] A conical combination of two entropy vectors realized by hypergraphs is realized by the hypergraph obtained by the disjoint union of the hypergraphs with the weights appropriately rescaled., and we have Walter and Witteveen (2021):
Theorem 1.
The entropy cone of hypergraph models is contained in the entropy cone of stabilizer states.
While in principle hypergraph models are not sufficient to solve the EQMIP completely at arbitrary ${\sf N}$ (see §VII), they at least provide a partial solution that will be sufficient for our purposes.
Refer to caption222200111122223333444455556666
Figure 1: A hypergraph giving an element of $\mathscr{R}_{6}$. The hyperedge (yellow) has weight 2, and edges (blue) have weight 1.
Refer to caption222211112222333344445555666677778888999900
Figure 2: A hypergraph giving an element of $\mathscr{R}_{9}$. The hyperedge (yellow) has weight 2, and edges (blue) have weight 1.
Refer to caption22222222222222222222222222223333444455556666222211110077778888
Figure 3: A hypergraph giving an element of $\mathscr{R}_{8}$. The two hyperedges (yellow and purple) have weight 2, and edges (blue) have weights either 1 (when unspecified) or 2.
VI Constructions for small ${\sf N}$
Having reviewed the necessary tools from hypergraph models, we now use this technology to construct inner bounds for $\overline{\Sigma}^{*}_{{\sf N}}$ for ${\sf N}\leq 4$. The logic will be as follows: we will construct hypergraphs that realize certain extreme rays of the SAC which, by 1, are automatically solutions to the EQMIP. Even without a full solution to the EQMIP, we will then show via particular coarse-grainings that we can obtain entropy vectors for a smaller number of parties, which give powerful inner bounds.
VI.1 Bounds for ${\sf N}=2$
This case it trivial, but we include it for completeness. The outer bound $\Sigma_{2}$ to $\overline{\Sigma}^{*}_{2}$ is the SAC, and its extreme rays are realized by Bell pairs. We then have
$\Delta_{2}^{2}=\overline{\Sigma}^{*}_{2}=\Sigma_{2}.$ (8)
VI.2 Bounds for ${\sf N}=3$
As we mentioned in §II, it was already shown in Pippenger (2003) that $\overline{\Sigma}^{*}_{3}=\Sigma_{3}$. This equivalence follows from the fact that the extreme rays of $\Sigma_{3}$ contain entropy vectors realized by Bell pairs, the 4-party GHZ state, and the 4-party absolutely maximally entangled state (also known as the 4-party perfect tensor) Helwig and Cui (2013). With the only exception of the 4-party GHZ, all these states also generate extreme rays of the 3-party SAC and are hence elements of $\mathscr{R}_{3}$. Their conical hull is $\Delta_{3}^{3}$, which coincides with the 3-party holographic entropy cone Bao et al. (2015).
To show that the sequence in Eq. (6) converges to $\overline{\Sigma}^{*}_{3}$, we need to consider $\Delta_{3}^{3+i}$ for $i\geq 1$. The sets $\mathscr{R}_{4}$ and $\mathscr{R}_{5}$ were already found in Hernández-Cuenca et al. (2019), and it is straightforward to verify that they imply $\Delta_{3}^{3}=\Delta_{3}^{4}=\Delta_{3}^{5}$. Hence, we need $i\geq 3$. It turns out that $i=3$ is sufficient, since one can verify that the entropy vector associated to the hypergraph in Figure 1 is in $\mathscr{R}_{6}$, and under the coarse-graining
$\{0,1,2,3,4,5,6\}\rightarrow\{0,1,1,2,2,3,3\},$ (9)
results in a hypergraph model of the GHZ state. In summary, we have
$\Delta_{3}^{3}=\Delta_{3}^{4}=\Delta_{3}^{5}\subset\Delta_{3}^{6}=\overline{\Sigma}^{*}_{3},$ (10)
which by Eq. (6) also implies $\Delta_{3}^{6+i}=\Delta_{3}^{6}$ for all $i\in\mathbb{N}$.
VI.3 Bounds for ${\sf N}=4$
As mentioned in §II, the QEC for ${\sf N}=4$ is not known. However, we will show that our construction, even with a relatively small value of ${\sf N}^{\prime}$, suffices to derive an inner bound that coincides with the subset of $\overline{\Sigma}^{*}_{4}$ of entropy vectors realized by stabilizer states. This subset, which we denote be $\overline{\Lambda}_{4}^{*}$, was shown in Linden et al. (2013) to coincide with its outer bound $\Lambda_{4}$, defined as the polyhedral cone specified by SA, SSA, and Ingleton’s inequality Ingleton (1971). Our goal is to show that there is some $i\in\mathbb{N}$ such that
$\Delta_{4}^{4+i}=\overline{\Lambda}_{4}^{*}.$ (11)
To see this, we can use a result from Bao et al. (2020a), which constructed hypergraph models realizing the entropy vectors that generate the extreme rays of $\Lambda_{4}$. It is then enough to show that each of these models (seven in total) can be obtained from coarse-graining another hypergraph involving ${\sf N}^{\prime}$ parties (for some ${\sf N}^{\prime}\geq 4$), whose corresponding entropy vector is an extreme ray of the ${\sf N}^{\prime}$-party SAC. Four out of these seven cases are trivial, since they correspond to elements of $\mathscr{R}_{4}$. For the remaining three non-trivial cases (which require $i>0$), we construct the hypergraphs shown in Figures 13. As one can verify, they all are all associated to extreme rays of the SAC for their corresponding number of parties. By an appropriate choice of coarse-grainings, we can reduce these hypergraphs to the ones labeled by 4, 6 and 7 in Figure 4 of Bao et al. (2020a) (up to a trivial permutation of the parties). Specifically, these coarse-grainings are
$\begin{split}\{0,1,2,3,4,5,6\}&\rightarrow\{0,1,1,2,2,3,4\}\\ \{0,1,2,3,4,5,6,7,8,9\}&\rightarrow\{0,1,1,2,2,3,3,4,4,0\}\\ \{0,1,2,3,4,5,6,7,8\}&\rightarrow\{0,1,1,2,2,3,3,4,4\}.\end{split}$
We then have
$\Delta_{4}^{4}=\Delta_{4}^{5}\subset\Delta_{4}^{6}\subset\overline{\Lambda}_{4}^{*}\subseteq\Delta_{4}^{9}\subseteq\overline{\Sigma}^{*}_{4}\subseteq\Sigma_{4},$ (12)
where the first equality follows trivially from the explicit knowledge of $\mathscr{R}_{4}$ and $\mathscr{R}_{5}$ Hernández-Cuenca et al. (2019), and the strict inclusion of $\Delta_{4}^{6}$ in $\overline{\Lambda}_{4}^{*}$ follows from the results of He et al. regarding $\mathscr{R}_{6}$. It is possible that $\mathscr{R}_{8}$, or even $\mathscr{R}_{7}$, might be enough to obtain $\overline{\Lambda}_{4}^{*}$, and we leave this for future work.
VII Discussion
We conclude by commenting on a few open questions regarding the strength of our bound, which we believe to be particularly interesting.
For the bounds that we have explicitly constructed in this letter, we relied on the hypergraph models of Bao et al. (2020a) and the result of Walter and Witteveen (2021). However, it was shown in Bao et al. (2020b) that for ${\sf N}\geq 5$ the inclusion of the entropy cone of hypergraphs in the cone of stabilizer states is strict. This implies that hypergraph models might not be sufficient to derive the most stringent possible bounds. On the other hand, we cannot rule out the possibility that hypergraph models do suffice to solve the EQMIP for arbitrary ${\sf N}$, in which case our bounds would not converge to the full QEC for any ${\sf N}\geq 4$. To resolve this, one direction is to look for elements in $\mathscr{R}_{{\sf N}}$ that can be realized by stabilizer states but violate the “hypergraph” inequality found in Bao et al. (2020b) at ${\sf N}=5$.
A similar limitation might also affect stabilizer states, since it is possible that for some value of ${\sf N}$, the solution to the EQMIP requires non-stabilizer quantum states. This question is related to the most stringent bound that can be realized at ${\sf N}=4$. Indeed, any $\Delta_{4}^{4+i}$ such that $\overline{\Lambda}_{4}^{*}\subset\Delta_{4}^{4+i}$ requires an element of $\mathscr{R}_{4+i}$ that cannot be realized by a stabilizer state, since it has to violate Ingleton’s inequality. Such violations have already been investigated for classical probability distributions Mao and Hassibi (2009), and another question for the future is to determine if for some ${\sf N}$ there exist extreme rays of the SAC that violate Ingleton’s inequality while being realizable by quantum states (at least approximately).
More generally, it seems conceivable that deep structural properties of the QEC might be extractable by exploring the behavior of the sequence of approximations given in Eq. (6). One possibility, already mentioned in §IV, is that the approximation does not improve beyond a certain point. This was indeed the case for ${\sf N}=3$, but as we have shown, in that case the approximation is exact, and perhaps this is not surprising given that $\overline{\Sigma}^{*}_{3}$ is polyhedral. On the other hand, it is possible that for larger ${\sf N}$, $\overline{\Sigma}^{*}_{{\sf N}}$ is not polyhedral, and it cannot be obtained as the conical hull of a finite set of vectors. Indeed, this is already known to be true for the subset of $\overline{\Sigma}^{*}_{{\sf N}}$ corresponding to classical probability distributions Matús (2007). Is polyhedrality related to the behavior of Eq. (6)? Does the sequence converge? If so, what is its limit, and how is it related to $\overline{\Sigma}^{*}_{{\sf N}}$? We leave these very interesting questions for future work.
Acknowledgments: T.H. is supported by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2817, the U.S. Department of Energy grant DE-SC0011632, and the Walter Burke Institute for Theoretical Physics. V.H. has been supported in part by the U.S. Department of Energy grant DE-SC0009999 and by funds from the University of California. M.R is supported by funds from the University of California.